# AsynEIO: Asynchronous Monocular Event-Inertial Odometry Using Gaussian Process Regression

### v1

@@ Abstract
Event cameras, when combined with inertial sensors, show significant potential for motion estimation in challenging scenarios, such as high-speed maneuvers and low-light environments. There are many methods for producing such estimations, but most boil down to a synchronous discrete-time fusion problem. However, the asynchronous nature of event cameras and their unique fusion mechanism with inertial sensors remain underexplored. In this paper, we introduce a monocular event-inertial odometry method called AsynEIO, designed to fuse asynchronous event and inertial data within a unified Gaussian Process (GP) regression framework. Our approach incorporates an event-driven frontend that tracks feature trajectories directly from raw event streams at a high temporal resolution. These tracked feature trajectories, along with various inertial factors, are integrated into the same GP regression framework to enable asynchronous fusion. With deriving analytical residual Jacobians and noise models, our method constructs a factor graph that is iteratively optimized and pruned using a sliding-window optimizer. Comparative assessments highlight the performance of different inertial fusion strategies, suggesting optimal choices for varying conditions. Experimental results on both public datasets and our own event-inertial sequences indicate that AsynEIO outperforms existing methods, especially in high-speed and low-illumination scenarios.
Index Terms:
event-inertial fusion, Gaussian process regression, motion estimation.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
The visual inertial odometry (VIO) technology is critical for autonomous robots to navigate in unknown environments. The VIO systems estimate motion states of robots by means of fusing Inertial Measurement Units (IMUs) and visual sensors. Generally, the Inertial Measurement Units (IMUs) directly measure angular velocities and linear accelerations while the visual sensors observe the bearing information with respect to environments. The complementarity of inertial and visual sensors makes the estimation problem more stable in both static and dynamic scenarios. Nevertheless, the conventional VIO systems are still prone to crash in several challenging scenarios, including high-speed movements, High-Dynamic-Range (HDR) environments, and sensors with incompatible frequencies. Specifically, high-speed movements and HDR environments will cause low-quality grayscale images for intensity cameras and long-period preintegration for IMUs. The long-period preintegration rapidly accumulates integration errors, meanwhile, low-quality images intend to deteriorate the visual frondend of VIO systems. Both of them result in a decline in estimation performance. In addition, discrete-time estimation and preintegration methods inherently have little support to fuse gyroscopes and accelerometers with incompatible frequencies or even asynchronous measurements. Therefore, it is imperative to develop a more robust VIO system to enhance estimation performance in challenging situations.
@@
Continuous-time estimation methods have been proposed to deal with foregoing drawbacks of event-based VIO systems. For these methods, a continuous-time trajectory is firstly defined with splines or kinematic assumptions, and then employed to build the measurement residuals. As the continuous-time trajectory have valid definitions at arbitrary timestamps, it naturally has the capacity of inferring the motion trajectory from totally asynchronous measurements and sensors. Among them, the Gaussian Process (GP) based methods have gained great interests of researchers for its clearer physical meaning. Although previous researchers have proposed to leverage GP-based methods to infer the SE(3) trajectory for event-based VIOs, existing methods still struggle to address the aforementioned challenges. With the introduction of GP-based methods, various inertial fusion schemes combining event cameras and IMUs have emerged. Thus, the interesting question is: which is the best fusion scheme for event-based VIO?
In this paper, we suggest an Asynchronous Event-Inertial Odometry (AsynEIO) system to improve the performance in challenging cases and comprehensively assess the reasonable inertial schemes. This study significantly builds on our previous conference publication [7] by developing a purely event-based frontend and introducing new inertial fusion schemes. The proposed AsynEIO utilizes a fully asynchronous fusion pipeline based on GP regression to improve the estimation performance in HDR and high-speed environments. Two main types of GP-based inertial fusion methods are integrated into the system to permit to fuse incompatible frequencies or asynchronous inertial measurements (i.e., GP-inertial Factor (GPIF) and GP-preintegration (GPP)), as depicted in Fig. 1. The GPIF is specifically designed for GP-based continuous-time methods to avoid the need for integrating raw measurements. We achieve this objective by leveraging a White Noise on Jerk (WNOJ) motion prior and investigating the underlying relationship between the differential SE(3) state variables and raw inertial measurements. Instead, the GPP infers data-driven GP models from real-time inertial measurements and generate a series of latent states to replace the motion prior in GPIF. Inspired by GPIF, we further extend the conventional discrete preintegration (Preint) to develop a new GP-based inertial factor, referred to as ExtPreint, where new residuals are directly incorporated into the GP trajectory. We fairly compare the performance of these inertial fusion methods under the unified event-inertial pipeline. Eventually, we conduct experiments on both public datasets and own-collected event-inertial sequences to demonstrate the advantages of the proposal with respect to the state-of-the-art. In summary, the main contributions of this article are as follows:
@@
    A fully asynchronous event-inertial odometry named as AsynEIO that fuses pure event streams with inertial measurements using a unified GP regression framework.
@@
    Comparative assessments of different inertial fusion schemes where four types of inertial schemes (i.e., GPIF, GPP, Preint, and ExtPreint) are compared to demonstrate the advantages of different fusion schemes.
@@
    Comprehensive experiments conducted on both public datasets and own-collected sequences to show the improvements of our proposal with respect to the state-of-the-art.
The remainder of this paper is organized as follows. Sec. II reviews the relevant literature. The GP regression on $SE(3)$ with WNOJ are detailed in Sec. III. The GPP and GPIF are introduced in Sec. IV and Sec, V, respectively. The proposed AsynEIO with different inertial factors is described in Sec. VI. Experimental assessments are carried out in Sec. LABEL:sec:experiments. Finally, conclusions are drawn in Sec. VII.
@@ II Related Work
Event-based visual-inertial odometry has gained significant research interest for challenging scenarios such as aggressive motion estimation as well as HDR scene. The most critical issue in this field is how to fuse data from the synchronous IMU data and the asynchronous event streams [5]. Zhu et al. [8] constructed a batch event packets aligning it by two Expectation-Maximization[9] steps and fused with pre-integration IMU in an EKF backend. Mahlknecht et al. [10] synchronized measurements from the EKLT tracker [11] using extrapolation when a fixed number of events is triggered. Then, these associated visual measurements was fused with IMU under a filter-based backend. However, the frontend computational complexity of these methods is very high. Rebecq et al. [12] detected corners on motion compensated event image and tracked through Lucas Kanade (LK) optical flow, and then fused the IMU-preintegrated measurements in a standard keyframe-based framework. Furthermore, this work was extended to the Ultimate-SLAM [4] by combing the standard frames. Lee et al. [13] proposed a 8-DOF warping model to fuse events and frames for accurate feature tracks. The residual from frontend optimization process was used to update an adaptive noise covariance. There is another trend of using the Time Surface (TS) [14] ,which stores the timestamp of the recent event at each pixel, as event representation to apply frame-based detection and tracking technique. [15] performed Arc* [16] for corner detection and LK tracking on TS with polarity. A tightly coupled graph-based optimization framework was designed to obtain estimation from the IMU and visual observations. PL-EVIO [17] combined the event-based line and point features and image-based point features to provide more information in human-made scene based on [15]. ESVO [18] proposed a first event-based stereo visual odometry by maximizing the spatio-temporal consistency using TS. Extended from [18], Junkai et al. [19] proposed a gyroscope measurements prior via pre-integration to circumvent the degeneracy issue. Kai et al. [20] utilized an adaptive decay-based TS to extract features and proposed a polarity-aware strategy to enhance the robustness, Then, these tracked features and IMU were fused in the MSCKF state estimator Overall, all these methods transform asynchronous event streams to synchronous data association and convert high rate IMU data into inter frame motion constrain through IMU-preintegration.
@@
Another category of works directly processed event streams asynchronously [21, 16, 22, 23] and applied continuous-time state estimation methods to fuse asynchronous data. Unlike estimating discrete states, continuous-time estimation methods assume a continuous trajectory and approximates it through a set of basis functions. Interpolating on continuous trajectories, it can relate measurements observed at arbitrary times like multi cameras system [24], Lidar odometry [25, 26, 27] and event camera [28, 29]. Mueggler et al. [30] firstly used cubic B-splines trajectory representation to directly integrate the information tracked by a given line-based maps. Furthermore, to estimate the absolute scale, they fused pre-integration IMU between adjacent control points and extended [30] to point-based maps in [31]. Lu et al. [32] proposed event-based visual-inertial velometer which leveraged a B-spline continuous-time formulation to fuse the heterogeneous measurements from a stereo event camera and IMU. Besides the B-spline method, GP regression models the system’s motion state using GP and construct a GP prior between adjacent discrete states, like White-Noise-on-Acceleration (WNOA) prior [33] and WNOJ prior [34]. Wang et al. [28] applied a motion-compensated RANSAC and WNOA motion prior to a stereo event-based VO pipeline. Comparing the above two continuous trajectory representations, Johnson et al. [35] demonstrated the accuracy and computation time of WNOJ prior GP-based method and spline-based method are similar. Note that there is no work to implement the WNOJ prior to event-based odometry.
@@
Recently, researchers have focused on designing new inertial fusion formulas specially for continuous-time estimation frameworks. Some of research works have proposed continuous-time formulations of IMU pre-integration [36, 37, 38]. These methods used inertial measurements to characterize the system’s continuous motion rather than direct prior like B-spline or GP-prior. [36] formulated a GP-based preintegration method for one-axis rotations and upsample the gyroscope data to compute the rotation preintegration when rotating on multiple axes. This strategy was employed in line features [39] and corner features [40] event-based VIO. IDOL [39] extracted event clusters of line segments and tracked lines between consecutive event windows. they leveraged [36] to associate asynchronous event measurements. However, this method shown a high computational cost for its full-batch optimization. Dai et al. [40] proposed a asynchronous tracking method relied on decaying event-based corners and adopted a similar fusing scheme as[39]. More recently, Gentil et al. [38] proposed a latent variable model to analytically integrate inertial measurements and implicitly induce a data-driven GP. This method formulated a unified GP-based integration to address the issue of continuous integration over SE(3) in [36]. Although their methods is specially designed for GP frameworks, they still integrate the inertial measurements rather than directly inferring on the raw measurements. Burnett et al. [41] introduced a velocity measurement model with integration in linear velocity under the WNOA motion prior [42]. Then, they further improve their methods with the Singer motion prior [27] where angular velocity and body-centric acceleration [43]. However, both of their methods accomplished poor preciseness in the accelerometer measurement model. Zheng et al. [25] developed a multi-LiDAR and IMU state estimator based on the GP-based trajectory. They separately defined the kinematic state on SO(3) and linear vector spaces rather than SE(3) and achieved a direct residual model on the raw measurements of IMUs. Chamorro et al. [44] investigated the event-IMU fusion strategies under Kalman filter which compared the conventional predict-with-IMU-correction-with-vision method and predict-with-model-correction-with-measurement. However, these solutions may face challenges in scenarios where straight lines are absent. Our previous work [7] investigated two different GP-based methods to fuse asynchronous visual measurements and IMU preintegration. Building on this foundation, we further extend our previous work to incorporate more inertial fusion methods and conduct comprehensive experiments to evaluate the performance of different schemes in fusing asynchronous measurements. In addition, our previous work relies on intensity images as auxiliary measurements for tracking event streams, limiting its performance in high-speed and HDR scenarios. In AsynEIO, we directly track pure event streams, achieving excellent adaptability in these challenging conditions.
III Sparse GP regression on $SE(3)$
@@ III-A Preliminaries and Kinematics
Define the transformation of the rigid body frame $b$ with respect to (w.r.t) the fixed inertial frame $i$ as a matrix $\bm{T}_{ib}$, where the pose $\bm{T}_{ib}$ forms the special Euclidean group, denoted $SE(3)$. Suppose the WNOJ prior to the kinematics of $\bm{T}_{ib}$ and apply the right multiply update of $SE(3)$. We have the differential equation as follows:
@@
is the generalized velocity, $\bm{s}$ is the generalized acceleration, and $\bm{Q}_{c}$ is the power spectral density matrix of the GP. Given the detailed rotation and translation components as
@@ III-B GP Regression on Local Linear Space
To linearize it, the local variable $\bm{\xi}$ in the tangent space of $\bm{T}_{k}$ can be given by
@@
where $\bm{\mathcal{J}}(\bm{\xi}_{k}(t))$ is the right Jacobian of $\bm{\xi}_{k}(t)$. Notice that $\bm{\mathcal{J}}\approx\bm{1}^{6\times 6}$ when $\bm{\xi}_{k}(t)$ is small enough. Further take the time derivative of (8), we can show the relationship between $\bm{\varpi}(t)$ and $\ddot{\bm{\xi}}_{k}(t)$:
$\ddot{\bm{\xi}}_{k}(t)=\bm{\mathcal{J}}^{-1}(\bm{\xi}_{k}(t))\dot{\bm{\varpi}}(t)+\frac{d}{dt}\bm{\mathcal{J}}^{-1}(\bm{\xi}_{k}(t))\bm{\varpi}(t).$ (9)
Apply the first-order approximation and substitute (8), we have
@@
where $\bm{\phi}\in\mathbb{R}^{3}$ is the rotation component and $\bm{\rho}\in\mathbb{R}^{3}$ is the translation component. Now, we can define the local linear kinematic state w.r.t $\bm{T}_{k}$ as

### v2

# IEEE Copyright Notice

@@ Abstract
Event cameras, when combined with inertial sensors, show significant potential for motion estimation in challenging scenarios, such as high-speed maneuvers and low-light environments. While numerous methods exist for producing such estimations, most boil down to solving a synchronous discrete-time fusion problem. However, the asynchronous nature of event cameras and their unique fusion mechanism with inertial sensors remain underexplored. In this paper, we introduce a monocular event-inertial odometry method called AsynEIO, designed to fuse asynchronous event and inertial data within a unified Gaussian Process (GP) regression framework. Our approach incorporates an event-driven front-end that tracks feature trajectories directly from raw event streams at a high temporal resolution. These tracked feature trajectories, along with various inertial factors, are integrated into the same GP regression framework to enable asynchronous fusion. With deriving analytical residual Jacobians and noise models, our method constructs a factor graph that is iteratively optimized and pruned using a sliding-window optimizer. Comparative assessments highlight the performance of different inertial fusion strategies, suggesting optimal choices for varying conditions. Experimental results on both public datasets and our own event-inertial sequences indicate that AsynEIO outperforms existing methods, especially in high-speed and low-illumination scenarios.
@@ I Introduction
The Visual Inertial Odometry (VIO) technology is critical for autonomous robots to navigate in unknown environments. The VIO systems estimate motion states of robots by means of fusing Inertial Measurement Units (IMUs) and visual sensors. Generally, the IMUs directly measure angular velocities and linear accelerations while the visual sensors observe the bearing information with respect to environments. The complementarity of inertial and visual sensors makes the estimation problem more stable in both static and dynamic scenarios. Nevertheless, the conventional VIO systems are still prone to crash in several challenging scenarios, including high-speed movements, High-Dynamic-Range (HDR) environments, and sensors with incompatible frequencies. Specifically, high-speed movements and HDR environments will cause low-quality grayscale images for intensity cameras and long-period preintegration for IMUs. The long-period preintegration rapidly accumulates integration errors, meanwhile, low-quality images intend to deteriorate the visual frondend of VIO systems. Both of them result in a decline in estimation performance. In addition, discrete-time estimation and preintegration methods inherently have little support to fuse gyroscopes and accelerometers with incompatible frequencies or even asynchronous measurements. Therefore, it is imperative to develop a more robust VIO system to enhance estimation performance in challenging situations.
@@
Continuous-time estimation methods have the potential to address foregoing drawbacks of event-based VIO systems. For these methods, a continuous-time trajectory is firstly defined with splines or kinematic assumptions, and then employed to build the measurement residuals. As the continuous-time trajectory have valid definitions at arbitrary timestamps, it naturally has the capacity of inferring the motion trajectory from totally asynchronous measurements and sensors. Among them, the Gaussian Process (GP) based methods have gained great interests of researchers for its clearer physical meaning. Although previous researchers have proposed to leverage GP-based methods to infer the $SE(3)$ trajectory for event-based VIOs, existing methods still struggle to address the aforementioned challenges. With the introduction of GP-based methods, various inertial fusion schemes combining event cameras and IMUs have emerged. Thus, the interesting question is: which is the best fusion scheme for event-based VIO?
In this paper, we suggest an Asynchronous Event-Inertial Odometry (AsynEIO) system to improve the performance in challenging cases and comprehensively assess the reasonable inertial schemes. This study significantly extends our previous conference publication [7] by developing a purely event-based front-end and introducing new inertial fusion schemes. The proposed AsynEIO utilizes a fully asynchronous fusion pipeline based on GP regression to improve the estimation performance in HDR and high-speed environments. Two main types of GP-based inertial fusion methods are integrated into the system to permit the fusion of incompatible frequencies or asynchronous inertial measurements (i.e., GP-Inertial Factor (GPIF) and GP-Preintegration (GPP)), as depicted in Fig. 1. The GPIF is specifically designed for GP-based continuous-time methods to avoid the necessity of integrating raw measurements. We achieve this objective by leveraging a White-Noise-on-Jerk (WNOJ) motion prior and investigating the underlying relationship between the differential $SE(3)$ state variables and raw inertial measurements. Instead, the GPP infers data-driven GP models from real-time inertial measurements and generates a series of latent states to replace the motion prior in GPIF. Inspired by the GPIF, we further extend the conventional discrete Preintegration (Preint) to develop a new GP-based inertial factor, termed ExtPreint, where new inertial residuals are directly incorporated into the GP trajectory. An additional variant named $\text{GPP}^{\ast}$ is developed to incorporate the original GPP with WNOJ priors. We fairly compare the performance of these inertial fusion methods under a unified event-inertial pipeline. Eventually, we conduct experiments on both public datasets and own-collected event-inertial sequences to demonstrate the advantages of the proposal with respect to the state-of-the-art. In summary, the main contributions of this article are as follows:
@@
    A fully asynchronous event-inertial odometry, named AsynEIO, where a asynchronously parallel, event-triggered front-end is presented to track feature trajectories from raw event streams in real-time.
@@
    Various asynchronous fusion methods are introduced to infer motion trajectories from totally asynchronous measurements in a unified GP-based framework.
@@
    Comprehensive evaluation experiments conducted on public datasets and own-collected sequences demonstrate that the proposed pipeline exhibits superior performance in high-speed scenarios. Informed by the comprehensive evaluation, practical guidance is provided for making reasonable choices of asynchronous fusion schemes under various conditions.
The remainder of this paper is organized as follows. Sec. II reviews the relevant literature. The GPIF and GPP are detailed in Sec. III and IV, respectively. The overall pipeline of AsynEIO is described in Sec. V. Experimental assessments are carried out in Sec. VI. The discussion of the results is presented in Sec. VII. Finally, conclusions are drawn in Sec. VIII.
@@ II Related Work
Event-based visual-inertial odometry has gained significant research interest for its estimation capabilities in challenging scenarios, such as aggressive maneuvering and HDR environments. A key issue in this field is how to temporally fuse data from the synchronous, high-rate IMU (e.g., 1 kHz) and the asynchronous event camera [5]. Zhu et al. [8] aligned batch event packets by two Expectation-Maximization[9] steps and fused them with IMU preintegration measurements using an Extended Kalman Filter (EKF). Mahlknecht et al. [10] synchronized measurements from an EKLT tracker [11] using extrapolation triggered by a fixed event count. Then, these associated visual measurements were fused with IMU data within a filter-based back-end. However, their front-end schemes incur significant computational overhead. Rebecq et al. [12] proposed to fuse motion compensated event frames and preintegration measurements using a standard keyframe-based framework, where the corner event features were tracked by the vanilla Lucas-Kanade (LK) optical flow. This work was further extended to Ultimate-SLAM [4] by incorporating standard intensity frames. Lee et al. [13] proposed an 8-DOF warping model to fuse events and frames for accurate feature tracking. The residual terms from the front-end optimization process were used to update an adaptive noise covariance. The Time Surface (TS) [14], which stores timestamps of recent events, is another viable option for event representation when applying frame-based detection and tracking techniques. For example, Guan et al. [15] adopted Arc* [16] for corner detection and implemented LK tracking on the TS. They designed a graph-based back-end to realize tightly-coupled sensor fusion. They further presented PL-EVIO [17], which employs multi-modal features (points and Lines) and sensor data (events and frames) to enhance performance in human-made environments. Zhou et al. [18] proposed the first stereo event odometry system that maximizes spatio-temporal consistency on the TS. On this basis, Junkai et al. [19] designed a gyroscope-based prior to mitigate the degeneracy issue. Tang et al. [20] suggested an adaptive decay-based TS for feature detecting, where a polarity-aware strategy was applied to enhance robustness, and leveraged the MSCKF state estimator for sensor fusion. Overall, these methods transform asynchronous event streams into synchronous data associations and convert high-rate IMU data into inter-frame motion constraints through IMU preintegration.
@@
Another category of works processed raw event streams asynchronously [21, 16, 22, 23] and applied continuous-time state estimation methods to fuse asynchronous data. As demonstrated by Cioffi et al. [24], continuous-time estimation methods are normally superior to its discrete counterpart when the sensors are not strictly time-synchronized. Unlike estimating discrete states, continuous-time methods normally assume a continuous trajectory and approximates it through a set of basis functions. By interpolating on continuous trajectories, these methods can relate measurements observed at arbitrary times, such as in multi-camera systems [25], LiDAR odometry [26, 27, 28] and event cameras [29, 30]. Mueggler et al. [31] first used cubic B-spline trajectory representation to directly integrate information tracked from line-based maps. They further extended their approach to point-based maps and fused preintegration measurements between adjacent control points to estimate absolute scale [32]. Lu et al. [33] proposed an event-inertial velometer which leveraged a B-spline continuous-time formulation to fuse the heterogeneous measurements from a stereo event camera and IMU. Besides B-spline-based methods, GP regression models the system’s motion state using GP and constructs a GP prior between adjacent discrete states, such as the White-Noise-on-Acceleration (WNOA) prior [34] and the WNOJ prior [35]. Wang et al. [29] applied motion-compensated RANSAC and a WNOA motion prior to a stereo event odometry pipeline. Comparing the two continuous trajectory representations mentioned above, Johnson et al. [36] demonstrated that the accuracy and computational cost of WNOJ prior GP-based methods and spline-based methods are similar. To the best of our knowledge, there is still a lack of work on GP-based event-inertial odometry using the WNOJ prior.
@@
Recently, researchers have focused on designing new inertial fusion formulas specifically for continuous-time estimation frameworks. Several studies have proposed continuous-time formulations for IMU preintegration [37, 38, 39]. These methods use inertial measurements to characterize the system’s continuous motion rather than relying on direct priors such as B-splines or GP priors. Le et al. [37, 38] formulated a gyroscope preintegration method, named LPM, where a GP-based latent model was adopted to upsample and infer more precise rotation preintegration. This strategy has been employed in event-based VIO systems to fuse IMU data with line features [40] and corner features [41]. IDOL [40] extracted event clusters of line segments and tracked line features between consecutive event windows. They leveraged LPM [37] to associate IMU data with asynchronous event measurements. However, this method shown a high computational cost due to its full-batch optimization. Dai et al. [41] proposed an asynchronous tracking method relying on decaying event-based corners and adopted a fusion scheme similar to that in [40]. More recently, Gentil et al. [39] proposed a GP-preintegration (i.e., GPP) method where the latent variable optimization model was applied to analytically infer preintegration measurements and implicitly induce local data-driven GP trajectories. They leveraged six independent GP trajectories to address continuous preintegration on $SE(3)$. Although their method is specifically designed for GP-based frameworks, it still requires redundant integration and optimization steps rather than directly operating on a GP-based $SE(3)$ trajectory using raw measurements. Burnett et al. [42] introduced continuous-time radar-inertial and LiDAR-inertial odometry, adopting a velocity observation model under the WNOA motion prior [43] by partially integrating inertial measurements. Then, they further improved their method by incorporating angular velocity and body-centric acceleration residuals [44] with the Singer motion prior [28]. However, they reported poor accuracy in their accelerometer measurement model. Zheng et al. [26] developed a multi-LiDAR and IMU state estimator using two independent GP trajectories on $SO(3)$ and $\mathbb{R}^{3}$. In contrast, we present a direct observation model (GPIF and its variant ExtPreint) of raw inertial measurements for the differential states of a unified $SE(3)$ GP trajectory. In addition, Chamorro et al. [45] investigated the event-inertial fusion strategies within a Kalman filter framework, comparing the conventional “predict-with-IMU-correct-with-vision” method against the “predict-with-model-correct-with-measurement” approach. However, these solutions may face challenges in environments where straight lines are scarce or absent. Our previous work [7] investigated two different GP-based methods (i.e., Preint and GPP) to fuse asynchronous visual measurements and inertial preintegration. Building on this foundation, we further incorporate more inertial fusion methods within a unified GP-based framework and conduct comprehensive experiments to evaluate their performance in fusing asynchronous measurements. Moreover, the previous work relied on intensity images as auxiliary measurements for tracking event streams, limiting its performance in high-speed and HDR scenarios. In AsynEIO, we propose an asynchronous front-end to enable direct processing of pure event streams, achieving excellent adaptability in these challenging conditions.
III GP-Inertial Factor
The GPIF is directly built upon the continuous-time $SE(3)$ GP trajectory with the WNOJ motion prior. We first define the system kinematics and provide the specific GP regression formulations on the given kinematic states. Then, the observation models relating the differential $SE(3)$ state to raw inertial measurements are presented along with their analytical Jacobians.
@@ III-A Preliminaries and Kinematics
Define the transformation of the rigid body frame $b$ with respect to (w.r.t) the fixed inertial frame $i$ as a matrix $\bm{T}_{ib}$, where the pose $\bm{T}_{ib}$ forms the special Euclidean group, denoted $SE(3)$. Suppose the WNOJ prior to the kinematics of $\bm{T}_{ib}$ and apply the right multiply update of $SE(3)$. The kinematic differential equation can be expressed as follows:
@@
is the generalized velocity, $\bm{s}$ is the generalized acceleration, and $\bm{Q}_{c}$ is the power spectral density matrix of the GP. The above equations are consistent with that in [35], except that we adopt a right-perturbation update convention. Given the detailed rotation and translation components as
@@ III-B GP Regression on Local Linear Space
The following derivations follow [28], except that we employ a right-perturbation. To linearize the nonlinear SDEs, the local variable $\bm{\xi}$ in the tangent space of $\bm{T}_{k}$ can be given by
@@
where $\bm{\mathcal{J}}(\bm{\xi}_{k}(t))$ is the right Jacobian of $\bm{\xi}_{k}(t)$. Notice that $\bm{\mathcal{J}}\approx\bm{1}^{6\times 6}$ when $\bm{\xi}_{k}(t)$ is small enough. Further take the time derivative of (8) and apply the first-order approximation, we can show the relationship between $\bm{\varpi}(t)$ and $\ddot{\bm{\xi}}_{k}(t)$:
@@
where $\bm{\phi}\in\mathbb{R}^{3}$ is the rotation component and $\bm{\rho}\in\mathbb{R}^{3}$ is the translation component. Now, the local linear kinematic state can be defined w.r.t $\bm{T}_{k}$ as